# Strongly magnetized accretion with low angular momentum produces a weak jet

### v1

@@ 3.1.2 The role of eruptions
Figure 4: Three-dimensional rendering of the temporal evolution of plasma-$\beta$ and magnetic field lines, plotted on a scale of $50r_{g}$ during the first magnetic flux eruption episode for $r_{b}=100r_{g}$ (top row, a-d) and $r_{b}=500r_{g}$ (bottom row, e-h). Time progresses from left to right. The origin of magnetic field lines is fixed at two regions: around the BH and at $\approx 25r_{g}$ to the right of the BH in the image plane.
We illustrate the distinct dynamics of eruption episodes in simulations with increasing $r_{b}$ in Figure 4. Similarly to Fig. 3, this figure shows a three-dimensional volume rendering of plasma-$\beta$ (in color) and magnetic field lines for $r_{b}=100r_{g}$ (top row, a-d) and $r_{b}=500r_{g}$ (bottom row, e-h), both for $a=0$. The scale of the plot is approximately $50r_{g}$. We fix the positions of the starting points of the field lines on a sphere around the BH, as well as at a distance of about $25r_{g}$ from the BH, along the trajectory of the first eruption.
@@
When the first magnetic flux eruption occurs, it distorts the magnetic field lines (b and f). Figure 4 shows that this distortion is much more dramatic for $r_{b}=500r_{g}$ than for $r_{b}=100r_{g}$ (compare panels g and h to c and d). Our interpretation is that the smaller $r_{b}$ case has an unphysically high magnetic pressure at $r\sim 50-100r_{g}$ (relative to a more realistic larger $r_{b}$ case) which suppresses the outward propagation of the eruption. This in turn suppresses how efficiently the eruption can drive turbulence in the ambient medium and reorganize the inflowing magnetic field. In the more realistic larger $r_{b}$ case (bottom row in Fig. 4) the BH-feedback driven turbulence much more dramatically changes the magnetic field structure on scales of $\sim 100r_{g}$. We show below that it can even reverse the sign of the accreted magnetic flux.
@@
Figure 6: Mean profiles (averaged over a spherical shell and over 5 snapshots over time of $1000r_{g}/c$) of density (a), density-weighted plasma-$\beta$ (b), and radial physical velocity (c). The time frames are: $[39-40]\times 10^{3}$ for $r_{b}=100r_{g}$ and $a=0.95$; $[39-40]\times 10^{3}$ for $r_{b}=500r_{g}$ and $a=0.95$; $[35-36]\times 10^{3}$ for $r_{b}=100r_{g}$ and $a=0$; $[100-101]\times 10^{3}$ for $r_{b}=500r_{g}$ and $a=0$.
We also demonstrate profiles of mean quantities of the accretion flow in Figure 6, which are averaged over spherical shells at each radius (where $\sigma<10$ to avoid the jet region) over 10 snapshots spanning $1000r_{g}/c$ during a period of steady $\phi_{\rm BH}$. These profiles illustrate density (a), density-weighted plasma-$\beta$ (b), and density-weighted radial velocity (c) computed in the tetrad basis (see Appendix B). Similar to Fig. 1, different colors represent different $r_{b}$, while thin and thick lines represent cases where $a=0$ and $a=0.95$, respectively. The density profiles are all shallower than the Bondi profile, $\rho\sim r^{-3/2}$, which is roughly consistent with $r^{-1}$ as reported in Ressler et al. (2021). The density is higher close to the BH for $r_{b}=500r_{g}$ because we begin each simulation with the identical $\rho_{0}=1$.
In all cases, we observe that the plasma-$\beta$ value drops below $1$ near the horizon as shown in panel b, where $\beta\equiv 1$ is represented by the dotted gray line. For the spinning scenarios, plasma-$\beta$ is generally lower, reaching unity, $\beta\approx 1$, within approximately $15r_{g}$. However, the profiles are similar across different Bondi radii for a given spin, implying a similar level of magnetic dominance near the BH. The radial velocities follow a similar dependence (c). We return to the evolution of the angular velocity in § 3.3.
@@ 3.1.3 Magnetic field evolution
Far from the BH, all profiles closely follow the initial values at all times, the $r_{b}=500$ magnetic flux is thus lower due to higher initial plasma-$\beta$. In the case of $r_{b}=100r_{g}$, we find that the profiles quickly reach a steady state within $20000r_{g}/c$ for both spin values (solid lines in a and b). In contrast, $r_{b}=500r_{g}$ continues to evolve with time due to the turbulent nature of the accreted magnetic field. The difference is more dramatic at $a=0.95$ (b), where $r_{b}=500r_{g}$ also shows sign changes in $\Phi$, indicating changes in the direction of magnetic flux.
@@
To further examine the time evolution of the magnetized turbulence, Figure 8 shows plasma-$\beta$ profiles for $r_{b}=500r_{g}$ in a similar way to Fig. 7; as before, panels (a) and (b) show non-spinning and spinning cases respectively. The averaging is performed over a spherical shell at each radius in the regions where $\sigma<10$ to avoid the jet. The figure is divided into two parts on the horizontal axis: the left side shows a region close to the BH on a logarithmic $r$, while the right side shows linear $r$ at larger radii.
Initially, both profiles follow $r^{3/2}$ dependence within the Bondi radius, consistent with a Bondi solution with a radial magnetic field. As they evolve, they reach $\beta<1$ inside $r\approx 10r_{g}$. For a non-spinning BH (a), the $\beta<1$ region extends smoothly to larger distances. For $a=0.95$, instead of a steady growth, we observe large fluctuations in the profile due to BH-driven turbulence. The magnetically dominated region of $\beta<1$ can extend up to $100r_{g}$ at times.
At large distances outside the Bondi radius, the $\beta$ profiles do not saturate – the low $\beta$ values continue to extend further from the BH in time. In the highly-spinning case (b), this extension occurs faster and reaches larger distances. We anticipate that regions of small $\beta$ will continue to extend outward at times later than shown in our simulations.
@@ 3.2 Parameter Dependence
In all simulations with larger Bondi radii, we observe the same main dynamical behavior – a decrease in the normalized magnetic flux saturation level (panels c and g), compared to the case of $r_{b}=100r_{g}$. Consequently, at $a=0.95$, the outflow energy efficiency decreases, leading to the absence of a powerful jet (panel h). Overall, $\dot{M}$, $\Phi_{\rm BH}$, and $\phi_{\rm BH}$ for $r_{b}=500r_{g}$ and $r_{b}=1000r_{g}$ reach similar values for both $a=0$ and $a=0.95$. At a higher Bondi radius of $5000r_{g}$, we find the saturated value of $\phi_{\rm BH}$ to be a little lower, compared to $r_{b}=1000r_{g}$.
We also test the sensitivity of our results on the initial $\beta_{0}$ at $r_{b}=500r_{g}$ and $a=0.95$. We conducted a run with $\beta_{0}=100$ (same as for $r_{b}=100r_{g}$) at our medium resolution, and two additional values at smaller resolution, $\beta_{0}=10$ and $\beta_{0}=10^{4}$ (third column, panels i-l). Our findings show that the main results converge for $\beta_{0}=10$ and higher. The magnetic flux is at a similar level (panels j and k). The energy outflow efficiency is small, indicating that no powerful jet appears.
@@
3.3 Late evolution and the impact of gas angular momentum
In this Section, we explore the dependence of our results on the final free parameter of our problem – angular momentum $l$. For the toroidal component of the 4-velocity, we define:
@@ 3.3 Late evolution and the impact of gas angular momentum
Figure 10: Time evolution of main flow characteristics (organized similarly to Fig. 1) for simulations with $r_{b}=500r_{g}$, $\beta_{0}=10^{2}$, $a=0.95$ and varying initial circularization radius, $r_{c}$, represented by different colors. Simulation that started with zero angular momentum in the flow, $r_{c}=0r_{g}$, is performed at a “medium” resolution, while all simulations with $r_{c}>0$ are performed at a “small” resolution.
@@
Initially, all simulations reach a high value of $\Phi_{\rm BH}$, similar to the $r_{c}=0$ case. The larger the $r_{c}$, the longer the period during which $\Phi_{\rm BH}$ remains at a high level. This elevated $\Phi_{\rm BH}$, as well as $\phi_{\rm BH}$, is similar across all $r_{c}$ values. This state is characterized by low $\dot{M}$, high efficiency $\eta$ exceeding $100\%$, and large oscillations in $\Phi_{\rm BH}$ – characteristics typical of the MAD state.
@@
Figure 11: Mean profiles of density (a), $\rho$-weighted plasma-$\beta$ (b), and radial (c) and angular velocities (d) at 4 distinct times for $r_{b}=500r_{g}$, $a=0.95$, and $r_{c}=5r_{g}$. Panels (e-h) show snapshots of density at these times in the poloidal plane. Angular velocity is normalized by the Keplerian velocity profile at $\theta=\pi/2$.
We find that the transitions between high and low $\Phi_{\rm BH}$ (and correspondingly high and low jet efficiency $\eta$) are driven by changes in the angular momentum distribution in the system. This is illustrated in Figure 11, where panels (a-d) show radial profiles of several quantities at four distinct times, and panels (e-h) show the corresponding snapshots of plasma density in the poloidal plane at these times. Panels (a-d) demonstrate the profiles of average density (a), $\rho$-weighted plasma-$\beta$ (b), and radial (c) and angular (d) physical velocities.
At $10\times 10^{3}r_{g}/c$, Fig. 10 demonstrates that $r_{c}=5r_{g}$ case is in a typical MAD state. The snapshot (Fig. 11e) shows that the flow is organized into a thick disk and a powerful jet is present. The angular velocity is positive and high. Later, at $75\times 10^{3}r_{g}/c$, $\Phi_{\rm BH}$ falls; the inflow becomes less organized, with the jet becoming weaker and tilted to the side (f). At this time, the angular momentum profile shows small or even negative values (d). The jet reappears at $95\times 10^{3}r_{g}/c$ (g), which again correlates with high angular momentum (d). Finally, the jet partially disappears at $120\times 10^{3}r_{g}/c$ (h), while the angular momentum magnitude decreases again (d).
@@
Figure 12: Time evolution of $\rho$-weighted profiles of physical components of the gas velocity, $\langle v^{\hat{\theta}}\rangle_{\rho}$, (a); and, $\langle v^{\hat{\phi}}\rangle_{\rho}$, (b) measured in LNRF, for the case of $r_{b}=500r_{g}$, $\beta_{0}=10^{3}$, $r_{c}=0$, $a=0.95$. The velocities are normalized by Keplerian velocity in LNRF at $\theta=\pi/2$.
When the BH is spinning, $a=0.95$, the angular momentum of the accreting material evolves over time even in the absence of initial angular momentum, $r_{c}=0$. In Figure 12, we show the evolution of the $\rho$-weighted profiles of physical velocities $v^{\hat{\theta}}$ (a) and $v^{\hat{\phi}}$ (b), normalized by the Keplerian velocity profile at $\theta=\pi/2$, computed in a locally non-rotating frame (see Appendix B). This simulation ($r_{b}=500r_{g}$ and $\beta_{0}=10^{3}$), was described in detail in § 3.1. The rotation is typically, but not always, aligned with the spin axis of the BH, e.g., when $v^{\hat{\phi}}\approx 0.2v_{\rm K}$ at $t\approx\times 10\times 10^{4}r_{g}/c$, $v^{\hat{\theta}}\approx-0.2v_{\rm K}$. Notably, we find that the powerful jet with $\eta\approx 100\%$ (as was shown in Fig. 1) is associated with an increased angular velocity aligned with the spin axis of the BH. Specifically, the jet appears at $t\approx 160\times 10^{3}r_{g}/c$, which corresponds to the high angular momentum state shown by the red lines with $v^{\hat{\phi}}\approx 0.5v_{\rm K}$ (b). As in our simulations with initial angular momentum, we expect the gas to eventually return to a low angular momentum phase, leading to the subsequent absence of a powerful jet. The long-term duty cycle of phases with strong jets is unclear.
@@ 4 Conclusion and discussion
Using ideal GRMHD simulations we study spherical accretion of plasma with low angular momentum and net magnetic flux. We find that this configuration leads to a magnetically dominated plasma close to the black hole, characterized by an average plasma-$\beta\leq 1$. However, despite an abundance of magnetic flux available in the system and the strong magnetization close to the black hole, some key features of MADs (Tchekhovskoy et al., 2011) are not present. Notably, the dimensionless magnetic flux on the BH horizon saturates below the nominal value of $\phi_{\rm BH}\approx 50$ typically observed in the MAD state. As a result, for a spinning BH, the system only produces a weak jet.
@@
We attribute this significant departure from the typical MAD state to the effect of turbulence driven in the ambient medium by horizon-scale flow dynamics. As magnetic flux at the horizon accumulates, it leads to magnetic reconnection events that drive bubbles of magnetized plasma out into the accretion flow. This turbulence suppresses the accretion of magnetic flux leading to a weaker jet. In comparison to the standard MAD state, we do not find large fluctuations in the magnetic flux $\Phi_{\rm BH}$ at the horizon, indicating that large-scale magnetic eruptions are not present in the saturated state. There are, however, continued small eruption episodes that drive turbulence in the ambient medium, as indicated by smaller-scale fluctuations in the horizon-scale magnetic flux (Fig. 1f-g). Interestingly, we find that $\phi_{\rm BH}$ and the jet power are roughly independent of the Bondi radius for $r_{b}=500r_{g}$ and above. The initial spherical GRMHD simulations of Ressler et al. (2020) focused on $r_{b}=200r_{g}$ and found larger jet power and horizon magnetic flux (as we do here for $r_{b}=100r_{g}$); we attribute this to the higher equatorial magnetic pressure at the (small) Bondi radius artificially suppressing the effect of BH-feedback driven turbulence on magnetic flux accumulation.
@@
In our interpretation, turbulence driven by episodic eruptions of magnetic flux accumulated on the black hole likely enhances the turbulent transport of magnetic flux outwards, which counteracts the laminar advection of magnetic flux inwards. Note that the inflow speed is much smaller than the free-fall velocity (Fig. 6c) so that modest levels of turbulence in the ambient medium can suppress the accumulation of magnetic flux on the black hole. It is unclear, however, why the net result is a dimensionless magnetic flux of $\phi\sim 10$ (for $a=0.95$) over a range of parameters. A mean-field model of the competition between radial advection and eruption-driven transport of magnetic flux could be valuable for understanding this better. [^2] Standard accretion disk models of flux transport (e.g., Lubow et al. 1994) are not applicable to the low angular momentum case considered here.
@@
To understand the transition between the initially non-rotating plasma simulated here and the rotating torus initial conditions that have been the focus of previous work (e.g., Tchekhovskoy et al. 2011), we carried out a series of simulations with increasing angular momentum, as parameterized by the circularization radius $r_{c}$ of the inflowing plasma (all in the regime of $r_{c}<r_{b}$). Finite angular momentum initially leads to more efficient horizon-scale magnetic flux accumulation and a stronger jet. We argue that rotation of the plasma efficiently shears out and de-correlates the magnetic eruptions, significantly suppressing the extent of the turbulence in the flow and thus enhancing the accumulation of magnetic flux onto the BH.
@@
One of the more striking results of our simulations is that the rotation of the plasma in the inner $\sim 10-100r_{g}$ undergoes secular evolution on long timescales, independent of whether the plasma is initially rotating or non-rotating. Simulations with and without initial rotation of the flow both experience phases of low and high angular momentum at later times.
@@
For initially non-rotating plasma accreting onto a spinning BH, a powerful jet, $\eta>100\%$, can emerge for short periods of time. These episodes are correlated with an increase in the angular momentum of the accreting plasma aligned with the spin axis of the BH. We argue that this angular momentum is due to a combination of angular momentum in the near horizon region (due to the spinning BH) transferred to large radii by magnetic torques and random angular momentum associated with the turbulence driven in the ambient medium.
@@
Our results indicate that both the angular momentum content and horizon-scale magnetic flux of radiatively inefficient accretion flows can evolve much more in time than previously appreciated. Moreover, we have shown that these two dynamical components are in fact intimately related: in our simulations, large dimensionless magnetic flux and powerful jets appear only when the plasma in the inner $\sim 10-100r_{g}$ has significant angular momentum (even though the plasma close to the BH is strongly magnetized with $\beta\lesssim 1$ in all cases we have studied). The angular momentum content of the plasma is not, however, just a property of the “initial condition” at large radii, but dynamically evolves due to strong magnetic torques and turbulence driven in the surrounding medium by jets and near-horizon magnetic eruptions. It is not at all clear what the long-timescale statistical steady state for any of our simulations actually is, and whether it is captured by the simulations we have carried out or whether the angular momentum, magnetic flux, and jet properties will continue to secularly evolve. It is also striking that the plasma continues to get more strongly magnetized at most radii as a function of time as indicated by the evolving $\beta$ profiles shown in Figure 8. More work is clearly needed to understand the very long-timescale evolution of these systems.

### v2

@@ 3.1.2 The role of eruptions
Figure 4: Three-dimensional rendering of the temporal evolution of plasma-$\beta$ (note that the opacity is decreasing as the plasma-$\beta$ is increasing) and magnetic field lines, plotted on a scale of $50r_{g}$ during the first magnetic flux eruption episode for $r_{b}=100r_{g}$ (top row, (a)-(d)) and $r_{b}=500r_{g}$ (bottom row, (e)-(h)). Time progresses from left to right. The origin of magnetic field lines for rendering is fixed at two regions: around the BH and at $\approx 25r_{g}$ to the right of the BH in the image plane.
We illustrate the distinct dynamics of eruption episodes in simulations with increasing $r_{b}$ in Figure 4. Similarly to Fig. 3, this figure shows a three-dimensional volume rendering of plasma-$\beta$ (in color) and magnetic field lines for $r_{b}=100r_{g}$ (top row, (a)-(d)) and $r_{b}=500r_{g}$ (bottom row, (e)-(h)), both for $a=0$. The scale of the plot is approximately $50r_{g}$. We fix the positions of the starting points of the field lines on a sphere around the BH, as well as at a distance of about $25r_{g}$ from the BH, along the trajectory of the first eruption.
@@
When the first magnetic flux eruption occurs, it distorts the magnetic field lines ((b) and (f)). Figure 4 shows that this distortion is much more dramatic for $r_{b}=500r_{g}$ than for $r_{b}=100r_{g}$ (compare panels (g) and (h) to (c) and (d)). Our interpretation is that the smaller $r_{b}$ case has an unphysically high magnetic pressure at $r\sim 50-100r_{g}$ (relative to a more realistic larger $r_{b}$ case) which suppresses the outward propagation of the eruption. This in turn suppresses how efficiently the eruption can drive turbulence in the ambient medium and reorganize the inflowing magnetic field. In the more realistic larger $r_{b}$ case (bottom row in Fig. 4) the BH-feedback-driven turbulence changes the magnetic field structure much more dramatically on scales of $\sim 100r_{g}$. We show below that it can even reverse the sign of the accreted magnetic flux.
@@
Figure 6: Mean profiles (averaged over a spherical shell and over five snapshots over time of $1000r_{g}/c$) of density (a), density-weighted plasma-$\beta$ (b), and radial physical velocity (c). The time frames are: $[39-40]\times 10^{3}$ for $r_{b}=100r_{g}$ and $a=0.95$; $[39-40]\times 10^{3}$ for $r_{b}=500r_{g}$ and $a=0.95$; $[35-36]\times 10^{3}$ for $r_{b}=100r_{g}$ and $a=0$; $[100-101]\times 10^{3}$ for $r_{b}=500r_{g}$ and $a=0$.
We also demonstrate profiles of mean quantities of the accretion flow in Figure 6, which are averaged over spherical shells at each radius (where $\sigma<10$ to avoid the jet region; throughout the paper, $\sigma=B^{2}/4\pi\rho c^{2}$ denotes the plasma’s magnetization) over 10 snapshots spanning $1000r_{g}/c$ during a period of steady $\phi_{\rm BH}$. These profiles illustrate density (a), density-weighted plasma-$\beta$ (b), and density-weighted radial velocity (c) computed on the tetrad basis (see Appendix B). Similar to Fig. 1, different colors represent different $r_{b}$, while thin and thick lines represent cases where $a=0$ and $a=0.95$, respectively. The density profiles are all shallower than the Bondi profile, $\rho\sim r^{-3/2}$, which is roughly consistent with $r^{-1}$ as reported in Ressler et al. (2021). The density is higher close to the BH for $r_{b}=500r_{g}$ because we begin each simulation with the identical $\rho_{0}=1$.
In all cases, we observe that the plasma-$\beta$ value drops below $1$ near the horizon as shown in panel (b), where $\beta\equiv 1$ is represented by the dotted gray line. For the spinning scenarios, plasma-$\beta$ is generally lower, reaching unity, $\beta\approx 1$, within approximately $15r_{g}$. However, the profiles are similar across different Bondi radii for a given spin, implying a similar level of magnetic dominance near the BH. The radial velocities follow a similar dependence (c). We return to the evolution of the angular velocity in § 3.3.
@@ 3.1.3 Magnetic field evolution
Far from the BH, all profiles closely follow the initial values at all times, the $r_{b}=500$ magnetic flux is thus lower due to higher initial plasma-$\beta$. In the case of $r_{b}=100r_{g}$, we find that the profiles quickly reach a steady state within $20000r_{g}/c$ for both spin values (solid lines in (a) and (b)). In contrast, $r_{b}=500r_{g}$ continues to evolve with time due to the turbulent nature of the accreted magnetic field. The difference is more dramatic at $a=0.95$ (b), where $r_{b}=500r_{g}$ also shows sign changes in $\Phi$, indicating changes in the direction of magnetic flux.
@@
To further examine the time evolution of the magnetized turbulence, Figure 8 shows plasma-$\beta$ profiles for $r_{b}=500r_{g}$ in a similar way to Fig. 7; as before, panels (a) and (b) show the nonspinning and spinning cases, respectively. The averaging is performed over a spherical shell at each radius in the regions where $\sigma<10$ to avoid the jet. The figure is divided into two parts on the horizontal axis: the left side shows a region close to the BH on a logarithmic $r$, while the right side shows linear $r$ at larger radii.
Initially, both profiles follow $r^{3/2}$ dependence within the Bondi radius, consistent with a Bondi solution with a radial magnetic field. As they evolve, they reach $\beta<1$ inside $r\approx 10r_{g}$. For a nonspinning BH (a), the $\beta<1$ region extends smoothly to larger distances. For $a=0.95$, instead of steady growth, we observe large fluctuations in the profile due to BH-driven turbulence. The magnetically dominated region of $\beta<1$ can extend up to $100r_{g}$ at times.
At large distances outside the Bondi radius, the $\beta$ profiles do not saturate – the low $\beta$ values continue to extend further from the BH in time. In the highly spinning case (b), this extension occurs faster and reaches larger distances. We anticipate that regions of small $\beta$ will continue to extend outward at times later than shown in our simulations.
@@ 3.2 Parameter Dependence
In all simulations with larger Bondi radii, we observe the same main dynamical behavior – a decrease in the normalized magnetic flux saturation level (panels (c) and (g)), compared to the case of $r_{b}=100r_{g}$. Consequently, at $a=0.95$, the outflow energy efficiency decreases, leading to the absence of a powerful jet (panel (h)). Overall, $\dot{M}$, $\Phi_{\rm BH}$, and $\phi_{\rm BH}$ for $r_{b}=500r_{g}$ and $r_{b}=1000r_{g}$ reach similar values for both $a=0$ and $a=0.95$. At a higher Bondi radius of $5000r_{g}$, we find the saturated value of $\phi_{\rm BH}$ to be a little lower, compared to $r_{b}=1000r_{g}$.
We also test the sensitivity of our results on the initial $\beta_{0}$ at $r_{b}=500r_{g}$ and $a=0.95$. We conducted a run with $\beta_{0}=100$ (same as for $r_{b}=100r_{g}$) at our medium resolution, and two additional values at a smaller resolution, $\beta_{0}=10$ and $\beta_{0}=10^{4}$ (third column, panels (i)-(l)). Our findings show that the main results converge for $\beta_{0}=10$ and higher. The magnetic flux is at a similar level (panels (j) and (k)). The energy outflow efficiency is small, indicating that no powerful jet appears.
@@
3.3 Late evolution and impact of gas angular momentum
In this section, we explore the dependence of our results on the final free parameter of our problem – angular momentum $l$. For the toroidal component of the four-velocity, we define:
@@ 3.3 Late evolution and impact of gas angular momentum
Figure 10: Time evolution of main flow characteristics (organized similarly to Fig. 1) for simulations with $r_{b}=500r_{g}$, $\beta_{0}=10^{2}$, $a=0.95$ and varying initial circularization radius, $r_{c}$, represented by different colors. The simulation that started with zero angular momentum in the flow, $r_{c}=0r_{g}$, is performed at a “medium” resolution, while all simulations with $r_{c}>0$ are performed at a “small” resolution.
@@
Initially, all simulations reach a high value of $\Phi_{\rm BH}$, similar to the $r_{c}=0$ case. The larger the $r_{c}$-value, the longer the period during which $\Phi_{\rm BH}$ remains at a high level. This elevated $\Phi_{\rm BH}$, as well as $\phi_{\rm BH}$, is similar across all $r_{c}$ values. This state is characterized by low $\dot{M}$, high efficiency $\eta$ exceeding $100\%$, and large oscillations in $\Phi_{\rm BH}$ – characteristics typical of the MAD state.
@@
Figure 11: Mean profiles of density (a), $\rho$-weighted plasma-$\beta$ (b), and radial (c) and angular velocities (d) at four distinct times for $r_{b}=500r_{g}$, $a=0.95$, and $r_{c}=5r_{g}$. Panels ((e)-(h)) show snapshots of density at these times in the poloidal plane; black lines represent the magnetic field lines. The angular velocity is normalized by the Keplerian velocity profile at $\theta=\pi/2$.
We find that the transitions between high and low $\Phi_{\rm BH}$ (and correspondingly high and low jet efficiency $\eta$) are driven by changes in the angular momentum distribution in the system. This is illustrated in Figure 11, where panels ((a)-(d)) show radial profiles of several quantities at four distinct times, and panels ((e)-(h)) show the corresponding snapshots of plasma density in the poloidal plane at these times. Panels ((a)-(d)) demonstrate the profiles of average density (a), $\rho$-weighted plasma-$\beta$ (b), and radial (c) and angular (d) physical velocities.
At $10\times 10^{3}r_{g}/c$, Fig. 10 demonstrates that $r_{c}=5r_{g}$ case is in a typical MAD state. The snapshot (Fig. 11e) shows that the flow is organized into a thick disk and a powerful jet is present. The angular velocity is positive and high. Later, at $75\times 10^{3}r_{g}/c$, $\Phi_{\rm BH}$ falls; the inflow becomes less organized, with the jet becoming weaker and tilted to the side (f). At this time, the angular momentum profile shows small positive values near the BH and negative values beyond $\approx 3r_{g}$ (d). The jet reappears at $95\times 10^{3}r_{g}/c$ (g), which again correlates with high angular momentum (d). Finally, the jet partially disappears at $120\times 10^{3}r_{g}/c$ (h), while the angular momentum magnitude decreases again (d).
@@
Figure 12: Time evolution of $\rho$-weighted profiles of physical components of the gas velocity, $\langle v^{\hat{\theta}}\rangle_{\rho}$, (a); and, $\langle v^{\hat{\phi}}\rangle_{\rho}$, (b) measured in the LNRF, for the case of $r_{b}=500r_{g}$, $\beta_{0}=10^{3}$, $r_{c}=0$, $a=0.95$. The velocities are normalized by the Keplerian velocity in the LNRF at $\theta=\pi/2$.
When the BH is spinning, $a=0.95$, the angular momentum of the accreting material evolves over time even in the absence of initial angular momentum, $r_{c}=0$. In Figure 12, we show the evolution of the $\rho$-weighted profiles of physical velocities $v^{\hat{\theta}}$ (a) and $v^{\hat{\phi}}$ (b), normalized by the Keplerian velocity profile at $\theta=\pi/2$, computed in a locally nonrotating frame (LNRF; see Appendix B). This simulation ($r_{b}=500r_{g}$ and $\beta_{0}=10^{3}$), was described in detail in § 3.1. The rotation is typically, but not always, aligned with the spin axis of the BH, e.g., when $v^{\hat{\phi}}\approx 0.2v_{\rm K}$ at $t\approx\times 10\times 10^{4}r_{g}/c$, $v^{\hat{\theta}}\approx-0.2v_{\rm K}$. Notably, we find that the powerful jet with $\eta\approx 100\%$ (as shown in Fig. 1) is associated with an increased angular velocity aligned with the spin axis of the BH. In particular, the jet appears at $t\approx 160\times 10^{3}r_{g}/c$, which corresponds to the high-angular-momentum state shown by the red lines with $v^{\hat{\phi}}\approx 0.5v_{\rm K}$ (b). As in our simulations with initial angular momentum, we expect the gas to eventually return to a low-angular-momentum phase, leading to the subsequent absence of a powerful jet. The long-term duty cycle of phases with strong jets is unclear.
@@ 4 Conclusion and discussion
Using ideal GRMHD simulations, we study spherical accretion of plasma with low angular momentum and net magnetic flux. We find that this configuration leads to a magnetically dominated plasma close to the black hole, characterized by an average plasma-$\beta\leq 1$. However, despite an abundance of magnetic flux available in the system and the strong magnetization close to the black hole, some key features of MADs (Tchekhovskoy et al., 2011) are not present. Notably, the dimensionless magnetic flux on the BH horizon saturates below the nominal value of $\phi_{\rm BH}\approx 50$ typically observed in the MAD state. As a result, for a spinning BH, the system only produces a weak jet.
@@
We attribute this significant departure from the typical MAD state to the effect of turbulence driven in the ambient medium by horizon-scale flow dynamics. As magnetic flux at the horizon accumulates, it leads to magnetic reconnection events that drive bubbles of magnetized plasma out into the accretion flow. This turbulence suppresses the accretion of magnetic flux leading to a weaker jet. In comparison to the standard MAD state, we do not find large fluctuations in the magnetic flux $\Phi_{\rm BH}$ at the horizon, indicating that large-scale magnetic eruptions are not present in the saturated state. There are, however, continued small eruption episodes that drive turbulence in the ambient medium, as indicated by smaller-scale fluctuations in the horizon-scale magnetic flux (Figures 1(f)-(g)). Interestingly, we find that $\phi_{\rm BH}$ and the jet power are roughly independent of the Bondi radius for $r_{b}=500r_{g}$ and above. The initial spherical GRMHD simulations of Ressler et al. (2020) focused on $r_{b}=200r_{g}$ and found larger jet power and horizon magnetic flux (as we do here for $r_{b}=100r_{g}$); we attribute this to the higher equatorial pressure at the (small) Bondi radius artificially suppressing the effect of BH-feedback-driven turbulence on magnetic flux accumulation.
@@
In our interpretation, turbulence driven by episodic eruptions of magnetic flux accumulated on the black hole likely enhances the turbulent transport of magnetic flux outward, which counteracts the laminar advection of magnetic flux inward. Note that the inflow speed is much smaller than the freefall velocity (Figure 6(c)) so modest levels of turbulence in the ambient medium can suppress the accumulation of magnetic flux on the black hole. It is unclear, however, why the net result is a dimensionless magnetic flux of $\phi\sim 10$ (for $a=0.95$) over a range of parameters. A mean-field model of the competition between radial advection and eruption-driven transport of magnetic flux could be valuable for understanding this better. [^3] Standard accretion disk models of flux transport (e.g., Lubow et al. 1994) are not applicable to the low-angular-momentum case considered here.
@@
To understand the transition between the initially nonrotating plasma simulated here and the rotating torus initial conditions that have been the focus of previous work (e.g., Tchekhovskoy et al. 2011), we carried out a series of simulations with increasing angular momentum, as parameterized by the circularization radius $r_{c}$ of the inflowing plasma (all in the regime of $r_{c}<r_{b}$). Finite angular momentum initially leads to more efficient horizon-scale magnetic flux accumulation and a stronger jet. We argue that rotation of the plasma efficiently shears out and decorrelates the magnetic eruptions, significantly suppressing the extent of the turbulence in the flow and thus enhancing the accumulation of magnetic flux onto the BH.
@@
One of the more striking results of our simulations is that the rotation of the plasma in the inner $\sim 10-100r_{g}$ undergoes secular evolution on long timescales, independent of whether the plasma is initially rotating or nonrotating. Simulations with and without initial rotation of the flow both experience phases of low and high angular momentum at later times.
@@
For initially nonrotating plasma accreting onto a spinning BH, a powerful jet, $\eta>100\%$, can emerge for short periods of time. These episodes are correlated with an increase in the angular momentum of the accreting plasma aligned with the spin axis of the BH. We argue that this angular momentum is due to a combination of angular momentum in the near-horizon region (due to the spinning BH) transferred to large radii by magnetic torques and random angular momentum associated with the turbulence driven in the ambient medium.
@@
Our results indicate that both the angular momentum content and horizon-scale magnetic flux of radiatively inefficient accretion flows can evolve much more in time than previously appreciated. Moreover, we have shown that these two dynamical components are in fact intimately related: in our simulations, large dimensionless magnetic flux and powerful jets appear only when the plasma in the inner $\sim 10-100r_{g}$ has significant angular momentum (even though the plasma close to the BH is strongly magnetized with $\beta\lesssim 1$ in all cases we have studied). The angular momentum content of the plasma is not, however, just a property of the “initial condition” at large radii, but dynamically evolves due to strong magnetic torques and turbulence driven in the surrounding medium by jets and near-horizon magnetic eruptions. It is not at all clear what the long-timescale statistical steady state for any of our simulations actually is, whether it is captured by the simulations we have carried out, or whether the angular momentum, magnetic flux, and jet properties will continue to secularly evolve. It is also striking that the plasma continues to get more strongly magnetized at most radii as a function of time as indicated by the evolving $\beta$ profiles shown in Figure 8. More work is clearly needed to understand the very long-timescale evolution of these systems.